An engineer designs cars to comfortably fit woman above the height of 157.5cm. The mean height of woman in the UK is 160.9 cm with a standard deviation of 6.0 cm. What proportion of women will be excluded from buying this car? (Assume the height of women is normally distributed)

Answer by stanbon(75887)

An engineer designs cars to comfortably fit woman above the height of 157.5cm. The mean height of woman in the UK is 160.9 cm with a standard deviation of 6.0 cm. What proportion of women will be excluded from buying this car? (Assume the height of women is normally distributed)
---------------------------
z(157.5) = (157.5-160.9)/6 = -0.5667
----------------------
P(x <=157.5) = P(z<=-0.5667) = 0.2855
-------------------------------------------
Cheers,
Stan H.